if the 2 sides of a triangle and the included angles r a=1+31/2 cm, b = 2 cm C = 600. solve the


triangle.

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear

a=1+31/2 cm, b = 2 cm C = 600

use cos formula

cos60 = (b2 + a2 -c2) /2ba

put the value

c= √6

now

  c/sinC = b /sinB

B=45

so A = 180-60-45

       =75

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin